Startup Night 10 - 20170911 UT

Transferred from Google Docs:

https://docs.google.com/document/d/1I3Epd407YOg7mzaXCG5rQgBELeVfzqy4Fk2gSmil9gk/edit

Tucson: Jenny, Olga, John Hill, Amali, Eckhart, Phil

Mountain: Geno, John Morris, Ilya, Klaus, Arto,

Summary: Variable and poor seeing plagued us. We attempted AGw transform during a brief stint of good seeing, but only had time to work out a few kinks and bugs before seeing blew up.

First light on both PepsiPol\x92s! Pepsi Guider workDid some work with LBC\x92s but /newdata FULL which put a kink in things. Data is still on the LBC PC\x92s. Handed off for Pepsi PFU work at the end of the night. Exercised those swing arms!

Night priorities outlined: https://docs.google.com/spreadsheets/d/1q3hQ04UESRPDJTzQn5SQ1EOXBOee_conayY7WJYgpRQ/edit#gid=0



Starting with LUCI Binocular Sky flats.

18:45 Geno opening

luci#.20170911

18:58 Track preset to blank field. Test 0001

K 10pt dither high luci1 & luci2 0002-00011

K 10pt dither low luci1 & 2 0012-0021

Ks 5pt dither low luci1 & 2 0022-0026

Ks 5pt dither high luci1&2 0027-0031

J test 0032

J low level luci1 0033 & 35-41 luci2 32-39

GEIRS sync error luci1 0034 is 2.51

J high level luci1 0042-0046 luci2 0040-0044

19:36 Seeing is ~1.5-2\x94 so handing over to LBTI until things settle

21:57 LBTI has completed their tasks

  1. Pointed to target, and adjusted M3 angle

  2. Checked for adequate room in M2 hexapod to correct coma on both sides + fill WFS pupils

  3. Checked out SX and tested old SX reconstructor post repairs on sky

  4. Send a command from adsec engineering gui simultaneous with an ICE command. IT 6358

To summarize, Happy people! SX is checked out and is working better than before the magnets were reattached.

21:57 Talked to Doug. Stone P still low and seeing not quite good enough to proceed with transform (1.2\x94 on source). Waiting to reevaluate in an hour. Moving on to Pepsi pol in the mean time.

22:11 PepsiPol preset. First light on both Pepsi Pol! DIMM Seeing 1.3-1.6\x94

23:11 Seeing showing some signs of improving

23:27 reconfiguring to LUCI - N30 Field Stop, H, lookup table ON

(Olga keeping log while Jenny running transform_collect on both sides)

23:42 starting script to get first off-axis exposures for dit, ndit estimates

List is Stone_P1_50.list (in ~lbto/idl/wfsc/transform/lists). Warning with Tip-Tilt request (oops, we had a preset to Stone_B1 instead\x85).

23:53:39 geirs error on LUCI2. Sync error. \x93GEIRS received respose with error: ERROR 1 Command \x91sync read\x92 returned errorcode = 1: (E_error=1) general error: \x91error: ERROR rad (E_ptimeout=21) timeout on OPTPCI interface\x92 sh: geirs_dropcaches: command not found.

Not fixed by reading GEIRS.

Restart LUCI2 GEIRS READOUT service via lmc. It took a long time (~1 min) to start again.

00:01 Now on Stone_P1 on-axis. Hiccup with copointing limit. Started transform_collect again which seemed to resolve this.

00:06 First LUCI image coming\x85 well, commit_readout injected two obsitems into the LUCI OEP, one for L1, one for L2, each with only a readout config which is 0s x 1 x 1, and then the script hung. What action to take? Used readout gui to submit pair of DIT x NDIT = 2.51 x 3 exposures, 47 and 48 for L1 and L2. But IDL script is still hung. Skyping Barry.

00:23 Can run a LUCI script and take an exposure. (Note that when we ran the LUCI script, the lookup table was turned off.) Restarting IDL. Run into copointing limit (Doug\x92s script must be using a star other than \x930\x94). Run transform_collect again. Success!

rm525-1 froze

luciN.20170911.nnnn.fits:

00:29 50,51 are skies on L1,L2

Stone_P1

11 50

10 50 51,50

4 50 52,51

2 50 53,52

12 50 54,53

9 50 xxx had to skip, guide probe out of range

1 50 55,54

7 50 xxx also bad, guide probe out of range

16 50 xxx also bad, gp out of range

6 50

00:46 Realized that on-axis star should be 0 and not 11. The last set is garbage. Editing list to have first line after name be \x930 50\x94

Stone_P1

sky 56,55

11 57,56

10 58,57 0.97\x94 on the guider, 1.53\x94 on the DIMM

4 59,58

2 60,59

12 61,60 star is off on L2 image

9 62,61

1 63,62 guide star on DX was jumping around a lot on previous on-axis collim (seeing 2\x94 on the DIMM and 1.8\x94 on the gider).

7 64,63

16 65, 64

6 66, 65

transform_luciN.log files written under /home/lbto/data/20170911/lfbg and rfbg.

01:10 Since seeing has deteriorated, we are stopping the collection of transform data and handing back to PEPSI for work on off-axis guiding. (Got off-axis guiding working, but one preset later AGw4 GP refused to move.)

1:14 Taking some luci darks during reconfig: 5 x 2.51sec

Luci2.20170911.0067-0071

Luci1 geirs sync error. Read detector then

Luci1.20170911.0069-73

01:23 Pepsi guider stuff: Off-axis guide star acquisition. Never succeeded in guiding off-axis with AGW4 as the guide probe was not working properly.

2:28 reconfi to LBC

2:38 preset with LBC to Feige 110, collimate and copointing

3:00 Feige 110 Uspec, r, F972, z

Lbcb.20170911.095812 -101440

Lbcr.20170911.095916-101416

3:24 Science field IC342. Collimate and pointing. Some cirrus moving in

3:39 IC342 z Uspec

3:56 refocus

4:03 IC342 long zUspec Seeing around 1\x94 on blue, 1.6 on red guider.

4:20 Images not displaying on RB_Science. Noticed that image size in /newdata of previous was undersized, then zero. Investigating. Handing off to pepsi. Newdata is full!! No errors in LBC log. On LBC PC\x92s images exist and are full size. Zero sized images not copied to repository but undersized ones were.

4:30 Geno reconfiguring for a record breaking 7th time. Pepsi is going to try and wrap up PFU work.



PEPSIPOL Setup

SX lab hotspot 279,247
DX lab hotspot 164,290
Installed those in PEPSIPOL_S.cfg

LBTI SX IE=+40 CA=+162 ---> SX_LBTI.20170907.ptmod
LBTI DX IE=+15 CA=+36 ---> DX_LBTI.20170907.ptmod

SX M3 ABS Tip=+600 Tilt=0 ---> LBTI SX IE=+40 CA=+128

This may have minor overlap with notes above.

05:10 Starting with PEPSIPOL gamEquA

SX IE=+29 CA=+14
DX IE=-26 CA=+17
using eyeball guiding and active optics

06:35

LUCI SX IE=+9 CA=+14
LUCI DX IE=-21 CA=+10

seeing got too bad

08:10 Switching back to PEPSIPOL

BS9188 has a close guide star

ACTIVE preset gs=0; DX collimated itself; cutout image is wrong; SX not on hotspot; +2000 Z7

08:37 ACQUIRE preset gs=1
ACQUIRE preset gs=2, 36.6,555.5 changing x-sign manually finds stars

08:42 XYFLIP to true/false

was +37.7,+56.8 for gs=2, but XYFLIP made it unreachable +37,-56
undo XYFLIP ---> false/false

08:52 M_XGP from +1 to -1 in oacontrol.conf
Now off-axis star appears on SX, but othing on DX.

08:55 ACQUIRE preset gs=0 to recenter pointing
PPOL SX IE=+15 CA=+6
PPOL DX IE=-33 CA=+14

ACTIVE preset to gs=2
SX almost collimating
DX probe failed to move

ACTIVE preset to gs=1
SX OK at -30.29,+19.93
DX failed -30.28,+19.93

09:13 GCSR stop; stop/start AGW4; probe wouldn't home.

09:18 switch to LBC

Comparing oacontrol.conf

(Use a grep like this: grep -e AGW -e M_YGP oacontrol.conf)

Variable 1 2 3 4
MXGP -1 -1 -1 -1
MYGP +1 +1 +1 +1
MXCCD -1 +1 -1 +1
MYCCD -1 +1 -1 +1
MXWFS +1 +1 +1 +1
MYWFS +1 +1 +1 +1
pcsmx -1 -1 -1 -1
pcsmy +1 +1 +1 +1

-- %USERSIG{BarryRothberg - 2017-09-11}%

Comments

 
Topic revision: r4 - 09 May 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback